OpenStudy (lgbasallote):

A baseball is hit so that it travels straight upward after being struck by the bat. A fan observes that it takes 3.00 s for the ball to reach its maximum height. Find the ball's initial velocity.

OpenStudy (lgbasallote):

we use \(x = V_{ox} t + \frac 12 at^2\) right?

OpenStudy (lgbasallote):

and a would be -9.8 m/s^2 right?

OpenStudy (lgbasallote):

wait doesn't sound right...

OpenStudy (shubhamsrg):

why dont you go for d(x)/dt = 0 => for maximum height a=-9.8 t=3 u= ? note that its the same thing as v-u = at v=0

OpenStudy (lgbasallote):

but what will you take the derivative of? there's no equation

OpenStudy (anonymous):

the derivative of position is velocity. so you get v = at

OpenStudy (anonymous):

woops, v = at + vo

OpenStudy (lgbasallote):

wait..you guys are using \[V = V_o + at\] right?

OpenStudy (lgbasallote):

so \[0 = V_o + (-9.8)(3)?\]

OpenStudy (anonymous):

yeah. thats the derivative of x = xo + vot + (1/2) a t^2

OpenStudy (lgbasallote):

lol never noticed that =))

OpenStudy (lgbasallote):

so is my equation right??

OpenStudy (anonymous):

yeah it is. then go back to the position equation, use that Vo, you already know the t and the a. Just get the x.

OpenStudy (lgbasallote):

so to find the maximum height i use x = vot + 1/2 at^2 right?

OpenStudy (anonymous):

exactly!

OpenStudy (lgbasallote):

great! thanks :D